OpenStudy (anonymous):

SOLVE FOR X: x^2 – 1 = x + 1

OpenStudy (anonymous):

x²-x-2 = 0 (x+1)(x-2)=0 x = -1 x = 2

OpenStudy (anonymous):

Thanks! I understand it now.